I’m one experiment 8.66g of NO is mixed with 5.03 g of O2 what is the limiting reactant


1
Expert's answer
2022-04-11T02:54:04-0400

2NO + O2 = 2NO2

n(NO) = m(NO) / Mr(NO) = 8.66 g / 14.01 g/mol = 0.618 mol

n(O2) = m(O2) / Mr(O2) = 5.03 g / 16.00 g/mol = 0.314 mol

2 mol NO react with 1 mol O2

NO is limiting reactant